# Estimating odds and log odds with guaranteed accuracy

### v2

@@ 1 Introduction
This paper deals with the problem of estimating $\gamma$ and $\phi$ from independent, identically distributed (i.i.d.) observations of a Bernoulli distribution with parameter $p$. Sequential estimators of $\gamma$ and $\phi$ are presented that achieve a specified level of accuracy, defined as relative or absolute MSE respectively. The sequential character of the estimators cannot be avoided, because a fixed sample size cannot satisfy these conditions for all $p$. The proposed estimators are based on inverse binomial sampling Haldane, 1945; Mikulski and Smith, 1976; Lehmann and Casella, 1998, chapter 2, and have a number of desirable features: they are simple to compute; they are unbiased; and their performance is close to optimal, in the sense that the relationship between accuracy and average sample size is only slightly worse than that given by Wolfowitz’s bound (Wolfowitz, 1947).
@@ 3 Estimation of log odds
For the first inequality in (28), consider the shaded rectangle in Figure 2(b), which extends from $x=k-1$ to $x=k$ and has height $1/k$, with $k\in\{r,\ldots,n-1\}$ arbitrary. Let $\alpha_{k}$ be chosen so that the intercept points of the curve $y=1/(x+1/2+\alpha_{k})$ with the vertical lines $x=k-1$ and $x=k$ are vertically equispaced from the horizontal line $y=1/k$, i.e. have heights $1/k+\beta_{k}$ and $1/k-\beta_{k}$ for some $\beta_{k}>0$. That is, $\alpha_{k}$ and $\beta_{k}$ satisfy the equation system
@@
The fraction with denominator $(2r-1)(r-1)$ in (50) is easily seen to be less than $1$ for $r\geq 2$, and thus (27) holds. ∎
@@
$\begin{split}\operatorname{Var}[\hat{\phi}]&<\left(\frac{1}{2}+\frac{1}{4r}\right)\left(\frac{1-p}{r-1+p}+\frac{p}{r-p}\right)+\frac{1}{2r-1}\\ &\quad+\frac{1}{4(r-1)}\left(\frac{p}{r-1+p}+\frac{1-p}{r-p}\right)\\ &=\frac{r-2p(1-p)}{2(r-1+p)(r-p)}+\frac{r-2p(1-p)}{4r(r-1+p)(r-p)}+\frac{1}{2r-1}\\ &+\frac{1}{4(r-1)}\,\frac{r-1+2p(1-p)}{(r-1+p)(r-p)}.\end{split}$ (57)
@@
By Theorem 3, $\hat{\phi}$ is an unbiased estimator of $\log(p/(1-p))$. Furthermore, choosing $r$ large enough guarantees, according to (54), an error variance as low as desired irrespective of $p$; at the cost of an increase in average sample size proportional to $r$, as given by (55). Thus, for example, to ensure an RMSE smaller than $0.2$ it suffices to use $r=27$ (which actually guarantees that the RMSE is less than $0.19701$). Reducing the target RMSE by a factor of $1/\sqrt{2}$, or the MSE by half, requires $r=52$.
@@
The lower bound (62) is compared in Figure 3 with efficiency values obtained by Monte Carlo simulation, using the same procedure as in Section 2. As can be seen, the bound is less tight in this case, compared with the bound (18) for $\eta(\hat{\gamma})$. This could be expected, because the expression of $\hat{\lambda}$, on which $\hat{\phi}$ is based, is more difficult to deal with than that of $\hat{\gamma}$ (in the proof of Theorem 2, $\operatorname{Var}[\hat{\lambda}]$ is expressed as the expected value of the square of a certain logarithm, which is then bounded by a simpler function). Still, the uniform bound (63) is greater than that for $\eta(\hat{\gamma})$, given by (19); and, as was the case for that estimator, it becomes tighter as $r$ increases.
@@ 4 Discussion and further work
The estimators are based on applying inverse binomial sampling separately with parameters $p$ and $1-p$, where the latter is obtained by swapping the role of successes and failures in the Bernoulli samples. This approach might appear wasteful, because information about both $p$ and $1-p$ could in principle be inferred using only one of the two sample sets. However, since the two probabilities are complementary, it is never the case that the two separate procedures require a large number of samples at the same time; and indeed performance turns out to be close to optimal, as stated.
On the other hand, using a single set of samples to infer $p/(1-p)$ or its logarithm would pose some problems. As follows from DeGroot (1959, theorem 4.1), there is no unbiased estimator of these parameters in inverse binomial sampling. For $(1-p)/p$ an unbiased estimator does exist; but then it is not possible to guarantee a relative MSE, because as $p$ grows the sample size decreases, when it should increase so as to reduce the MSE. By similar arguments, an unbiased estimator of $\log((1-p)/p)$ exists but is not useful either. For fixed-size sampling there are no unbiased estimators of any of these quantities (DeGroot, 1959, section 4); and in any case, with a fixed sample size it is not possible to guarantee a desired relative MSE for odds, or a desired MSE for log odds. Moreover, regardless of the rule used for determining sample size, using a single sample set can make the error analysis more difficult. Namely, if the estimator is expressed as the product or sum of two estimators (as has been done in Sections 2 and 3), those two estimators are not statistically independent, because they are based on the same observations.

### v3

@@ 1 Introduction
This paper deals with the problem of estimating $\gamma$ and $\phi$ from independent, identically distributed (i.i.d.) observations of a Bernoulli distribution with parameter $p$. Sequential estimators of $\gamma$ and $\phi$ are presented that achieve a specified level of accuracy, defined as relative or absolute MSE respectively. The sequential character of the estimators cannot be avoided, because a fixed sample size cannot satisfy these conditions for all $p$. The proposed estimators are based on inverse binomial sampling Haldane, 1945; Mikulski and Smith, 1976; Lehmann and Casella, 1998, chapter 2, and have a number of desirable features, in addition to guaranteeing a target accuracy: they are simple to compute; they are unbiased; and their performance is close to optimal, in the sense that the relationship between accuracy and average sample size is only slightly worse than that given by Wolfowitz’s bound (Wolfowitz, 1947).
@@ 3 Estimation of log odds
For the first inequality in (28), consider the shaded rectangle in Figure 2(b), which extends from $x=k-1$ to $x=k$ and has height $1/k$, with $k\in\{r,\ldots,n-1\}$ arbitrary. Let $\alpha_{k}>0$ be chosen so that the intercept points of the curve $y=1/(x+1/2+\alpha_{k})$ with the vertical lines $x=k-1$ and $x=k$ are vertically equispaced from the horizontal line $y=1/k$, i.e. have heights $1/k+\beta_{k}$ and $1/k-\beta_{k}$ for some $\beta_{k}>0$. That is, $\alpha_{k}$ and $\beta_{k}$ satisfy the equation system
@@
The fraction with denominator $(2r-1)(r-1)$ in (50) is easily seen to be less than $1$ for $r\geq 2$, $p\in(0,1)$, and thus (27) holds. ∎
@@
$\begin{split}\operatorname{Var}[\hat{\phi}]&<\left(\frac{1}{2}+\frac{1}{4r}\right)\left(\frac{1-p}{r-1+p}+\frac{p}{r-p}\right)+\frac{1}{2r-1}\\ &\quad+\frac{1}{4(r-1)}\left(\frac{p}{r-1+p}+\frac{1-p}{r-p}\right)\\ &=\frac{r-2p(1-p)}{2(r-1+p)(r-p)}+\frac{r-2p(1-p)}{4r(r-1+p)(r-p)}+\frac{1}{2r-1}\\ &\quad+\frac{r-1+2p(1-p)}{4(r-1)(r-1+p)(r-p)}.\end{split}$ (57)
@@
By Theorem 3, $\hat{\phi}$ is an unbiased estimator of $\log(p/(1-p))$. Furthermore, choosing $r$ large enough guarantees, according to (54), an error variance as low as desired irrespective of $p$; at the cost of an increase in average sample size proportional to $r$, as given by (55). Thus, for example, to ensure an RMSE smaller than $0.2$ it suffices to use $r=27$ (which actually guarantees that the RMSE is less than $0.1971$). Reducing the target RMSE by a factor of $1/\sqrt{2}$, or the MSE by half, requires $r=52$.
@@
The lower bound (62) is compared in Figure 3 with efficiency values obtained by Monte Carlo simulation, using the same procedure as in Section 2. As can be seen, the bound is less tight in this case, compared with the bound (18) for $\eta(\hat{\gamma})$. This could be expected, because the expression of $\hat{\lambda}$, on which $\hat{\phi}$ is based, is more difficult to deal with than that of $\hat{\gamma}$ (in the proof of Theorem 2, $\operatorname{Var}[\hat{\lambda}]$ is expressed as the expected value of the square of a certain logarithm, which is then bounded by a simpler function). Still, the uniform bound (63) is greater than the corresponding one for $\hat{\gamma}$, given by (19); and, as was the case for that estimator, it becomes tighter as $r$ increases.
@@ 4 Discussion and further work
Each of the two estimators is based on applying inverse binomial sampling separately with parameters $p$ and $1-p$, where the latter is obtained by swapping the role of successes and failures in the Bernoulli samples. This approach might appear wasteful, because information about both $p$ and $1-p$ could in principle be inferred using only one of the two sample sets. However, since the two probabilities are complementary, it is never the case that the two separate procedures require a large number of samples at the same time; and indeed performance turns out to be close to optimal, as stated.
On the other hand, using a single set of samples to infer $p/(1-p)$ or its logarithm would pose some problems. As follows from DeGroot (1959, theorem 4.1), there is no unbiased estimator of these parameters in inverse binomial sampling. For $(1-p)/p$ an unbiased estimator does exist; but then it is not possible to guarantee a relative MSE, because as $p$ grows the sample size decreases, when it should increase so as to reduce the MSE. By similar arguments, an unbiased estimator of $\log((1-p)/p)$ exists but is not useful either. For fixed-size sampling there are no unbiased estimators of any of these quantities (DeGroot, 1959, section 4); and in any case, with a fixed sample size it is not possible to guarantee a desired relative MSE for odds, or a desired MSE for log odds. Moreover, regardless of the rule used for deciding the sample size, using a single sample set can make the error analysis more difficult. Namely, if the estimator is expressed as the product or sum of two estimators (as has been done in Sections 2 and 3), those two estimators are not statistically independent, because they are based on the same observations.